OpenStudy (anonymous):

If f is continuous and (integral of f(x)dx from 0 to 4 is = 10), find the integral of (f(2x))dx from 0 to 2.

OpenStudy (anonymous):

What I did: for first integral: F(4)-F(0)=10 therefore, F(0)=F(4)-10 for second integral: F(2(2))-F(2(0)) = F(4) - F(4) + 10 =10 But the answer is 5. Where did I go wrong? Any help would be appreciated :)

OpenStudy (experimentx):

divide by 2